The total number of hours, measured in units of 100 hours, that a family runs a vacuum cleaner over a period of one year is a continuous random variable X that has the density function

1
Expert's answer
2022-06-14T11:17:43-0400
"f(x) = \begin{cases}
 x, &0<x<1 \\
 2-x, &1\le x<2 \\
0, & elsewhere
\end{cases}"

100 hours is taken as one unit of "X."

a) Find the probability that over a period of one year, a family runs their vacuum cleaner less than 120 hours


"P(X<1.2)=\displaystyle\int_{0}^{1}xdx+\displaystyle\int_{1}^{1.2}(2-x)dx"

"=[\dfrac{x^2}{2}]\begin{matrix}
 1\\
0
\end{matrix}+[2x-\dfrac{x^2}{2}]\begin{matrix}
 1.2\\
1
\end{matrix}"

"=0.5-0+2.4-0.72-2+0.5=0.68"

b) Find the probability that over a period of one year, a family runs their vacuum cleaner between 50 and 100 hours.


"P(0.5<X<1)=\displaystyle\int_{0.5}^{1}xdx"

"=[\dfrac{x^2}{2}]\begin{matrix}
 1\\
0.5
\end{matrix}=0.5-0.125=0.375"